# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-C3 Perception
Human visual attention and cognition provide a lot of information on how humans experience virtual spaces; however, the research in these fields does not directly answer how humans perceive these spaces. Considering perception engineering and its combining aspect of engineering and human physiology [269], many researchers have focused on the perception aspect of VR using eye movements as an important source of information. While some of the findings on human perception overlap with those on cognition and visual attention, it is important to address the findings individually and tackle the issues altogether to design more immersive and usable virtual spaces. In this part of the paper, we provide a detailed overview of the research that covers human perception of VR spaces through eye tracking.
Depth perception is a well-studied topic for human perception in VR. For instance, Katzakis et al. [270] investigated human depth perception during a 3D painting task in both foveal and peripheral regions. The study’s findings revealed that the task accuracy depended on the target’s location rather than the starting point, and highly accurate 3D painting tasks can be accomplished in virtual spaces when targets are within arm-reachable range. Nonetheless, the authors observed that overestimation of depth occurred in all cases, but it was more significant in the cases starting from the periphery and terminating in the foveal region. Further, Arefin et al. [271] revealed the relationship of perceptual depth level with eye vergence angle and interpupillary distance in VR, stating that these features could reveal the changes in perceptual depth.
Research on human perception also goes beyond depth perception. Serrano et al. [272] systematically investigated the human perception of continuity in VR movies, which have different cinematographic requirements from conventional 2D movies. The authors used gaze information and eye-tracking data with the videos and revealed new metrics such as scanpath errors describing visual attention, and offered design insights for VR content creation. Additionally, MacQuarrie and Steed [273] examined HMD users’ perception capacity to comprehend the eye-gaze direction of virtual volumetric characters in varying display resolutions, virtual character positions, head rotations, and gaze directions. The authors indicated that perceptual accuracy is considerably position-dependent and influenced by the direction of view.
@@
Identifying the perceptual limits of users is essential for VR applications that consist of reorientation and repositioning, which often require discrete or continuous scene changes. However, inserting the scene changes imperceptibly during the VR experience is essential. Optimizing the amount and timing of these changes according to users’ perceptual limits ensures that users’ virtual experience is not disturbed by the inserted changes. Researchers have examined VR users’ perceptual limits and abilities to have a detailed understanding of this and provide them with tailored VR solutions. For example, Shin et al. [274] examined the effect of unobtrusive and prompt virtual object movements on users’ body posture. The authors provided design insights while assessing users’ perceptual limits for unnoticeable position changes of the virtual objects during eye blinks. Langbehn et al. [275] investigated conscious or unconscious blinks that suppress human visual perception, which can be exploited to design motion algorithms in VR. The authors examined the upper limits of unnoticed changes during eye blinks and demonstrated that users could not perceive translational changes of 4-9 cm and rotations of $2^{\circ}-5^{\circ}$. Furthermore, they also stated that a 50% improvement can be observed in curvature gain, which is defined as $\frac{1}{r}$ where $r$ corresponds to the radius of a circular path in the real world while moving in a straight line in VR.
In addition to the blinks, Keyvanara and Allison [276] evaluated the human limits during saccadic behaviors and what effects and manipulations are imperceptible during saccadic suppressions that occur during constant horizontal and vertical camera movements by utilizing a Bayesian procedure. The authors reported that sudden horizontal camera movements, which were additionally applied to the continuously moving virtual camera on the vertical axis, were less observable. In another work, Keyvanara and Allison [277] examined the visual sensitivity of VR users against changes in a 3D scene during saccadic suppression. The authors’ findings revealed that certain image transformations, such as rotations along the roll axis, were more noticeable by users during horizontal saccades. Similarly, Bolte and Lappe [278] proposed using saccadic suppression to include rotations and translations into the scene. The authors modified two eye movement classification algorithms, which were initially proposed by Behrens et al. [279] and Niemenlehto [280] to meet the time and reliability requirements of online saccade detection. The authors investigated the limitations of human perception during detected saccadic periods and found that users could not perceive positional changes up to 50 cm in the direction of their eye gaze and rotational modifications up to $5^{\circ}$.
@@
Sense of realism is related to the visual characteristics of the 3D scenes we explore, the ecological validity of the simulations, and how we perceive those 3D spaces physiologically. As human visual attention drives eyes and we perceive the scenes accordingly, eye tracking analysis paves the way for understanding and estimating the realism level of virtual environments. This approach also allows the possibility of developing techniques for enhancing the perceived realism of VR environments, considering the perceptual limits of the human visual system.
@@
To this end, the effects of light sources on human vision, such as temporal eye adaptation, perceptual glare, visual acuity reduction, and scotopic color vision, can be used to enhance the realism level in virtual scenes. Luidolt et al. [281] utilized gaze direction and pupil size as means of lighting effect adjustment to produce more realistic low-light scenes that are tailored to the users’ vision. The authors revealed that the effect of the light in virtual scenes is highly subjective. Perception and reaction differences between virtual- and real-world tasks also provide insights to create realistic virtual environments. In this regard, eye-gaze activity is a crucial indicator for understanding human behavior in real and virtual worlds [282, 283].
@@
To design realistic crowd simulations, Berton et al. [282] investigated gaze behavior during crowd walking activity. The authors analyzed the differences in gazing activity in real and virtual simulations while there were varying crowd levels during virtual walking simulations. As the crowd level increased, participants’ eye movements became more restricted, scanning a narrower portion of the street. Furthermore, the authors noted that users directed their attention to the individuals in front of them. The authors also provided design recommendations, such as using a constant number of neighboring agents and positioning them by considering collision risk in virtual crowd simulations. In a different study, Berton et al. [283] exploited gaze data to compare collision avoidance behaviors in the real world and several VR environments. Their findings revealed that collision avoidance behaviors are similar across different settings. A work investigating the existence of the stare-in-the-crowd effect, which refers to the tendency of individuals to detect and observe the gaze directed to them, was designed by Raimbaud et al. [284]. By identifying differences in gaze characteristics, the authors revealed a stare-in-the-crowd effect. Additionally, the authors observed a negative correlation between the dwell time associated with this effect and social anxiety scores while providing insights to achieve realistic interactive environments with virtual avatars. Considering the similarities and differences between the real and virtual worlds, Gupta et al. [285] examined the autobiographical memory (AM) that is often considered as one of the vital factors in human perception. The authors revealed the effect of autobiographical memory on physiological cues like eye gaze, pupil diameter, and electrodermal activity.
@@
Avatars are essential components of virtual spaces and reality due to being a proxy for social interactions within virtual environments. They also have different impacts on perceived realism. To this end, eye-tracking methods to generate realistic avatars raised some attention. For example, Borland et al. [286] augmented a virtual self-avatar with eye movement animations to achieve a more comprehensive virtual embodiment and improve self-recognition. The authors compared representative animations of real eye movements with simulated eye movements without requiring eye-tracking hardware. The authors reported that use of eye movement animations increases the subjective sense of self-identification.
@@
Bergström et al. [287] evaluated the plausibility of virtual musical performance with changing environmental conditions and gaze-based attributes of audiences and musicians. The authors found that the gaze behavior of the virtual agents and distractions that do not comply with the nature of the environment had substantial effects. Ma and Pan [288] recently presented a framework to generate self-avatars with facial expressions and assessed the psychological effects of realistic avatars. In their study, participants found the facial expressions of cartoon-like avatars to be more controllable than those of realistic-looking avatars. Additionally, the authors observed that participants had a higher sense of body ownership in the first trial, regardless of the type of avatar used.
@@
Eyes, faces, and heads are important components of virtual avatars, and a handful of works have explicitly focused on simulating these components. For instance, Le et al. [289] introduced a method to automatically simulate head movement, eye gaze, and eyelid movement based on speech input. Gaussian Mixture Models and Nonlinear Dynamic Canonical Correlation Analysis were employed to simulate head movements and gaze behaviors. The authors applied non-negative linear regression for intentional blinks, while they generated unconscious blinks using a log-normal distribution. The results showed superior performance compared to state-of-the-art algorithms in generating head and eye motions. Thies et al. [290] introduced FaceVR, an image-based face synthesis approach that includes an eye-tracking algorithm based on monocular videos, aiming to render more realistic outputs on stereo displays. Ladwig et al. [291] proposed a low-cost solution to construct a person’s face wearing VR glasses utilizing GANs to produce face images. Additionally, the authors created an eye-tracking method that gives cues on iris position and gaze direction. Similarly, Song et al. [292] presented a CNN-based 3D face-eye reconstruction technique to construct a 3D personalized avatar with eye movements. The proposed algorithm can construct personalized avatars with the VR users’ facial expressions and eye animations. Another CNN-based technique introduced by Olszewski et al. [293] supports speech animation and emotional expressions, using mouth and eye-tracking cameras for HMDs. The proposed method outperforms state-of-the-art techniques regarding animation fidelity without requiring individual calibration. Further, Zhao et al. [294] proposed a face reconstruction algorithm using a personalized 3D head model along with a colorization algorithm for near-infrared eye images without causing the red-eye effect, which is defined as color distortion in the irises.
@@
While virtual spaces and setups provide users with many exciting opportunities, VR HMDs may lead to several types of physical discomfort, which are often called VR sickness or cybersickness, and are closely related to human perception. Eye tracking has often been used to understand the causes and severity of cybersickness and to develop methods to mitigate or prevent cybersickness. For example, Islam et al. [295] proposed a cybersickness severity prediction algorithm utilizing built-in HMD sensors, including eye and head trackers. This algorithm relies on deep fusion network models, and it was evaluated based on the data collected from 30 participants during a VR video game. The authors demonstrated the capability of their models by incorporating eye- and head-tracking features with video stimuli, and achieved an accuracy of up to 87.77%. In another work, Islam et al. [296] introduced a deep fusion network based on Deep Temporal Convolutional Networks (DeepTCN) fusing physiological, head-tracking, and eye-tracking data. DeepTCN could predict cybersickness 60 seconds in advance with a 0.49 mean-squared error (MSE) on a scale between 0 and 10. Similarly, Lee et al. [297] employed eye movement features to predict cybersickness severity along with disparity and optical flow maps that respectively represent depth and movement in the image. The authors’ 3D CNN model achieved better precision rates when it included eye-tracking features compared to Padmanaban et al. [298].
Cybersickness can occur in all types of virtual experiences; however, virtual experiences that require motion are more likely to cause this problem due to the mismatch between the movements in real and virtual worlds. For instance, virtual locomotion is one of the leading causes of VR sicknesses, such as nausea and dizziness, due to one-sided movement in the virtual world. Zayer et al. [299] provided a survey of virtual locomotion techniques that enhance the overall VR experience by mitigating motion sickness under the following categories: walking-based, steering-based, selection-based, and manipulation-based methods. The authors evaluated the strengths and weaknesses of these techniques to provide comprehensive guidance for the community. It is stated that motion sickness occurring during locomotion is due to the visual-vestibular conflict, and FOV constraints are deployed to mitigate VR discomfort by restricting users’ FOV. Adhanom et al. [300] presented a gaze position-based foveated FOV restriction method to improve existing techniques that rely on head gaze to mitigate cybersickness. These existing techniques are limited to cases where the head and eye gaze are aligned. The authors showed that their proposed method allows users to experience a wider visual scan field while preserving the same level of VR sickness and noticeability. When the relationship of cybersickness with locomotion techniques is considered, Qian and Teather [301] found that the technique that utilizes eye movements for input modality yields a high cybersickness due to the lack of head tracking.
@@
As users with different expertise and skills tend to have different visual perception patterns, the applications of eye-tracked VR have been developed with leaps and bounds in the realm of expertise training and assessment. Driving training is of particular interest among all VR training applications. For instance, Lang et al. [302] studied the improper driving habits of 50 users in a driving simulation with eye-tracking-enabled VR headsets and helped them improve their driving skills by designing synthesized personalized training routes considering their perceptual habits. Their user study conveyed that the 10 participants trained in a customized VR setup outperformed the others who were trained by conventional methods, with respect to response time in emergency situations, training persistence, and an evaluation score based on inappropriate driving actions (e.g., not signaling before a turn). In another domain, Orlosky et al. [303] applied eye tracking to measure and classify the English language understanding of 16 users in VR using features like pupil diameters and eye movements. The authors reached a 75% prediction accuracy for easy and medium words with SVMs. When they extended the analyses to include words that are categorized as hard, the prediction accuracies were up to 62%.
The applications in the industrial context are also important, considering these applications are used in everyday life. To this end, Burova et al. [304] developed an elevator maintenance simulation in eye-tracked VR to facilitate industrial AR prototypes, which are challenging to build due to safety concerns in the real world. The authors recorded several types of behavioral data from users, including eye behaviors, and enabled gaze data visualization through scanpaths and heatmaps during training playback in VR. Based on their survey, which included 12 elevator maintenance experts, the authors indicated that the domain experts hold constructive opinions regarding eye tracking and gaze visualizations for industrial training. Additionally, Gisler et al. [305] examined the relationship between training success and human behaviors in sanitary apprenticeship tasks. The authors combined gaze positions, head movements, and attention durations on the focused objects by using statistical summaries to predict the training success of users. Experiments including 48 sanitary apprentices in an industrial training task achieved a 10-20% improvement in predicting users’ training success compared to a baseline model. In addition, Harris et al. [306] examined skill training for searching and tracking in virtual police rooms. The authors designed virtual experiments in which officers were supposed to search and strive to gather evidence for an investigation while also considering various criminal activities. They analyzed the perceptual-cognitive skills of 54 participants in multiple training scenarios by taking saccadic eye movements into account. The results indicated that training users for visual search expertise in VR is possible.
@@
Previous research has also utilized eye tracking and VR to study vision impairment and ocular examination, as people with vision problems perceive the world differently. Like other subfields discussed in this paper, in VR, it is possible to have very controlled programmed experiments compared to in-the-wild settings to study such phenomena, which is why researchers have attempted to design ecologically valid simulations and study vision impairment issues in VR. For example, Krösl et al. [307] simulated cataract patients’ vision in a virtual environment to better understand the perception of people with cataracts. The authors used eye tracking for these gaze-dependent symptoms because cataracts lead to special visual effects at the lens center and periphery.
@@
More related to ocular examination using eye tracking in VR, Kim et al. [308] transplanted the Developmental Eye Movement test to VR HMDs, which is a clinical eye test widely used to determine abnormalities in visual function and to assess ocular motor skills. Subsequently, Kim et al. [309] designed the King-Devick test, which is a standard measurement for the assessment of saccadic eye movement and dynamic visual acuity, using HMDs both in VR and AR. Hotta et al. [310] also designed a gaze-based ocular examination for visual field defects in a virtual environment using fixation and saccadic features. Compared to conventional tests that take over 30 minutes, the proposed method can be conducted in five minutes while retaining sufficient accuracy and improving reliability. In addition, medical VR applications can go beyond the setups that require HMDs. For instance, Kübler et al. [311] analyzed patients’ eye and head movements with homonymous visual field defects (HVFD) in a virtual driving simulation. The virtual environment included a VR cabin, while a head-mounted eye tracker tracked participants’ eyes. The authors’ findings confirmed their hypothesis that a particular share of HVFD patients could strengthen their viewing behaviors with increased visual scanning to improve their driving skills. Similar to earlier sections, Table VIII presents an overview of the works on human perception.
@@
TABLE VIII: Overview of the papers related to human perception.
Paper # Participants Purpose Eye-tracking Features [270] 18 Depth perception analysis in 3D painting gaze [271] 24 Depth perception analysis ipd, eva [272] 49 Continuity analysis in VR movies gaze [273] 37 Perceptual limit analysis gaze [274] 27 Perceptual limit analysis blnk [275] 32 Perceptual limit analysis for redirected walking blnk [276] 36 Perceptual limit analysis sac [277] 10 Perceptual limit analysis sac [278] 13 Perceptual limit analysis sac [281] 5 Perceptual realism analysis in low-light 3D scenes gaze, pup [282] 21 Perception analysis in crowd walking gaze, fix [283] 17 Perception analysis in collision avoidance task gaze, fix [284] 30 Perception analysis in psychology fix [285] 20 Perception analysis gaze, pup, blink [286] 12 Perceived realism analysis for virtual avatars gaze [287] 20 Perceived realism analysis gaze [288] 18 Perceived realism analysis for virtual avatars gaze [289] 20 Perceived realism analysis for virtual avatars gaze [290] 18 Face & eye synthesis for virtual avatars N/A [291] 5 Face & eye synthesis for virtual avatars N/A [292] 7 Face & eye synthesis for virtual avatars N/A [293] 310 Face & eye synthesis for virtual avatars N/A [294] 3 Face & eye synthesis for virtual avatars N/A [295] 30 VR sickness prediction gaze, pup, cvgn [296] 30 VR sickness prediction gaze, pup, cvgn, blnk [299] N/A VR sickness mitigation N/A [297] 96 VR sickness prediction saliency [300] 22 VR sickness mitigation gaze [302] 50 Skill assessment and training for driving gaze [303] 16 Skill assessment and training in education pup, sac, fix [304] 12 Skill assessment and training for elevator maintenance fix, scanpath, saliency [305] 48 Skill assessment and training for sanitary apprentice gaze [306] 54 Skill assessment and training for police room search task gaze, sac, fix, entropy [307] 21 Vision simulation for patients with cataract gaze, pup [308] 39 Visual acuity assessment for ocular examination eye dominance, ipd, cvgn, phr [309] 30 Visual acuity assessment for ocular examination sac [310] 2 Visual field analysis for ocular examination sac, fix [311] 14 Visual field analysis in driving sac, fix, gaze

### v3

@@ IV-C3 Perception
Human visual attention and cognition provide a lot of information on how humans experience virtual spaces; however, the research in these fields does not directly answer how humans perceive them. Combining engineering and human physiology, researchers have also focused on the perception of VR, utilizing eye movements. While some findings on human perception overlap with those on cognition and visual attention, it is essential to address these findings individually and tackle the issues collectively to design immersive and usable virtual spaces. In this section of the paper, we provide a detailed overview of the research that examines human perception of VR using eye tracking.
Depth perception is a well-studied topic for human perception in VR. For instance, [239] investigated human depth perception during a 3D painting task in foveal and peripheral regions, revealing that task accuracy depended on the target’s location rather than the starting point. Highly accurate 3D painting tasks can be accomplished in virtual spaces when targets are arm-reachable. Nonetheless, the authors observed that overestimation of depth, and this overestimation was more significant in the cases starting from the periphery and terminating in the foveal region. Furthermore, [240] revealed the relationship between perceptual depth level and eye vergence angle and interpupillary distance in VR, stating that these features could reveal changes in perceptual depth.
Research on human perception also goes beyond depth perception. [241] systematically investigated the human perception of continuity in VR movies, which have different cinematographic requirements from conventional 2D movies. The authors utilized eye-tracking data in conjunction with the videos and revealed new metrics, such as scanpath errors, that describe visual attention, and provided design insights for VR content creation. Additionally, [242] examined HMD users’ perception capacity to comprehend the gaze direction of virtual volumetric characters in varying display resolutions, virtual character positions, head rotations, and gaze directions. The authors indicated that perceptual accuracy is considerably position-dependent and influenced by the direction of view.
@@
Identifying users’ perceptual limits is essential for VR applications that consist of reorientation and repositioning, which often require discrete or continuous scene changes. However, inserting the scene changes imperceptibly during the experience is essential. Optimizing the amount and timing of these changes according to users’ perceptual limits ensures that users’ virtual experience is not disturbed by these changes. To this end, researchers have examined the perceptual limits and abilities of users, particularly to provide them with tailored VR solutions. For example, [243] examined the effect of unobtrusive and prompt virtual object movements on users’ body posture. The authors provided design insights while assessing users’ perceptual limits for unnoticeable position changes of the virtual objects during eye blinks. [244] investigated conscious or unconscious blinks that suppress human visual perception, and demonstrated that users could not perceive translational changes of 4-9 cm and rotations of $2^{\circ}-5^{\circ}$. Furthermore, they also stated that a 50% improvement can be observed in curvature gain, which is defined as $\frac{1}{r}$ where $r$ corresponds to the radius of a circular path in the real world while moving straight in VR.
In addition to the blinks, [245] evaluated the human limits during saccadic behaviors and what effects and manipulations are imperceptible during saccadic suppressions that occur during constant horizontal and vertical camera movements by utilizing a Bayesian procedure. The authors reported that sudden horizontal camera movements, which were additionally applied to the continuously moving virtual camera on the vertical axis, were less observable. In another work, [246] examined the visual sensitivity of VR users against changes in a 3D scene during saccadic suppression. Findings revealed that certain image transformations, such as rotations along the roll axis, were more noticeable by users during horizontal saccades. Similarly, [247] proposed using saccadic suppression to include rotations and translations into the scene. The authors investigated the limitations of human perception during detected saccadic periods and found that users could not perceive positional changes up to 50 cm in the direction of their gaze and rotational modifications up to $5^{\circ}$.
@@
Sense of realism is related to the visual characteristics of the 3D scenes we explore, the ecological validity of the simulations, and how we perceive those 3D spaces physiologically. As human visual attention drives our eyes and we perceive scenes accordingly, eye-tracking analysis paves the way for understanding and estimating the realism level of virtual environments. This approach also enables the development of techniques to enhance the perceived realism of VR environments, taking into account the perceptual limitations of the human visual system.
@@
To this end, the effects of light sources on human vision, such as temporal eye adaptation, perceptual glare, visual acuity reduction, and scotopic color vision, can be used to enhance the realism level in virtual scenes. [248] utilized gaze direction and pupil size as means of lighting effect adjustment to produce more realistic low-light scenes that are tailored to the users’ vision. The authors revealed that the effect of the light in virtual scenes is highly subjective. Perception and reaction differences between virtual- and real-world tasks also provide insights into creating realistic virtual environments, in which gaze activity can be considered crucial for understanding human behavior in real and virtual worlds [249, 250].
@@
To design realistic crowd simulations, [249] investigated gaze behavior during crowd walking activity. The authors analyzed the differences in gazing activity in real and virtual simulations while there were varying crowd levels during virtual walking simulations. As the crowd level increased, participants’ eye movements became more restricted, scanning a narrower portion of the street. Furthermore, the authors noted that users directed their attention to the individuals in front of them and provided design recommendations, such as using a constant number of neighboring agents and positioning them by considering collision risks. In a different study, [250] exploited gaze data to compare collision avoidance behaviors in the real world and VR environments. Findings revealed that collision avoidance behaviors are similar across different settings. A work investigating the existence of the stare-in-the-crowd effect, which refers to the tendency of individuals to detect and observe the gaze directed at them, validated this effect, as well as a negative correlation between the dwell time associated with this effect and social anxiety scores, while providing insights to achieve realistic interactive environments with virtual avatars [251]. Considering the similarities and differences between the real and virtual worlds, [252] examined autobiographical memory (AM), which is often considered a vital factor in human perception. The authors revealed the effect of AM on physiological cues like eye gaze, pupil diameter, and electrodermal activity.
@@
Avatars are essential components of virtual spaces and reality, serving as proxies for social interactions. They also have different impacts on perceived realism. To this end, eye-tracking methods to generate realistic avatars raised some attention. For example, [253] augmented a virtual self-avatar with eye movement animations to achieve a more comprehensive virtual embodiment and improve self-recognition. The authors compared representative animations of real eye movements with simulated eye movements without requiring eye-tracking hardware and reported that the use of eye movement animations increases the subjective sense of self-identification.
@@
Considering the gaze behaviors of virtual agents, [254] evaluated the plausibility of virtual musical performance with changing environmental conditions and gaze-based attributes of audiences and musicians. The authors found that the gaze behavior of the virtual agents and distractions that do not comply with the nature of the environment had substantial effects. Furthermore, [255] recently presented a framework to generate self-avatars with facial expressions and assessed the psychological effects of realistic avatars. Participants found the facial expressions of cartoon-like avatars to be more controllable than those of realistic-looking avatars. Additionally, the authors observed that participants had a higher sense of body ownership in the first trial, regardless of the avatar type.
@@
Eyes, faces, and heads are essential components of virtual avatars, and a handful of works have explicitly focused on simulating those. For instance, [256] introduced a method, based on Gaussian mixture models and nonlinear dynamic canonical correlation analysis, and linear regression, to simulate head movement, eye gaze, and eyelid movement based on speech input. The results showed superior performance compared to state-of-the-art algorithms in generating head and eye motions. [257] introduced FaceVR, an image-based face synthesis approach that includes an eye-tracking algorithm based on monocular videos, aiming to render more realistic outputs on stereo displays. [258] proposed a low-cost solution to construct a person’s face wearing VR glasses, utilizing GANs to produce face images. Additionally, the authors created an eye-tracking method that gives cues on iris position and gaze direction. Similarly, [259] presented a CNN-based 3D face-eye reconstruction technique for constructing a personalized 3D avatar with eye movements. The proposed algorithm can construct personalized avatars with the VR users’ facial expressions and eye animations. Another CNN-based technique [260] supports speech animation and emotional expressions, using mouth and eye-tracking cameras for HMDs. The proposed approach outperforms state-of-the-art techniques in animation fidelity without requiring individual calibration. Furthermore, [261] proposed a face reconstruction algorithm using a personalized 3D head model along with a colorization algorithm for near-infrared eye images without causing the red-eye effect, which is defined as color distortion in the irises.
@@
While VR offers numerous exciting opportunities, HMDs may lead to several types of physical discomfort, often referred to as VR sickness or cybersickness, which are closely related to human perception. Eye tracking has been frequently employed to understand the causes and severity of cybersickness, as well as to develop methods for mitigating or preventing it. For example, [262] proposed a cybersickness severity prediction algorithm utilizing built-in HMD sensors, including eye and head trackers. This algorithm relies on deep fusion network models, and through the evaluations with a VR video game, the authors demonstrated the capability of their models by incorporating eye- and head-tracking features with video stimuli, and achieved an accuracy of up to 87.77%. In another work, [263] introduced a deep fusion network based on Deep Temporal Convolutional Networks (DeepTCN) fusing physiological, head-tracking, and eye-tracking data. DeepTCN could predict cybersickness 60 seconds in advance with a 0.49 mean-squared error (MSE) on a scale between 0 and 10. Similarly, [264] employed eye movement features to predict cybersickness severity along with disparity and optical flow maps that respectively represent depth and movement in the image. The authors’ 3D CNN model achieved better precision rates when it included eye-tracking features compared to previous work [265].
Cybersickness may occur in any virtual experience, but is more common in motion-based ones due to mismatched real and virtual movements. For instance, virtual locomotion is one of the leading causes of VR sicknesses, such as nausea and dizziness, due to one-sided movement in the virtual world. According to [266], sickness occurring during locomotion is due to the visual-vestibular conflict, and FOV constraints are deployed to mitigate VR discomfort by restricting users’ FOV. [267] presented a gaze position-based foveated FOV restriction method to improve existing techniques that rely on head gaze to mitigate cybersickness. These existing techniques are limited to cases where the head and eye gaze are aligned. The authors demonstrated that their proposed method enables users to experience a wider visual scan field while maintaining the same level of VR sickness and noticeability. Furthermore, considering the relationship between cybersickness and locomotion techniques, [268] found that the method utilizing eye movements for input modality yields high cybersickness due to the lack of head tracking.
@@
As users with different expertise and skills tend to have varying visual perception patterns, the applications of eye-tracked VR have advanced significantly in the realm of expertise training and assessment. Driving training is of particular interest among the applications. For instance, [269] studied improper driving habits in a driving simulation with eye-tracking-enabled VR headsets and helped participants improve their driving skills by designing personalized training routes that considered their perceptual habits. Participants who were trained in a customized VR setup outperformed those trained by conventional methods, with respect to response time in emergency situations, training persistence, and an evaluation score based on inappropriate driving actions (e.g., not signaling before a turn). In another domain, [270] utilized eye tracking to measure and classify English language understanding in VR using features such as pupil diameters and eye movements, and reached 75% and 62% prediction accuracies for easy and medium words, and for hard words, respectively.
Industrial applications are also important, as they are used in everyday life. To this end, [271] developed an elevator maintenance simulation in eye-tracked VR to facilitate industrial AR prototypes, which are challenging to build due to safety concerns in the real world. The authors recorded several types of behavioral data from users, including eye behaviors, and enabled gaze data visualization through scanpaths and heatmaps during training playback in VR. Based on their survey, the domain experts hold constructive opinions regarding eye tracking and gaze visualizations for industrial training. Additionally, [272] examined the relationship between training success and human behaviors in sanitary apprenticeship tasks. The authors combined gaze positions, head movements, and attention durations on the focused objects by using statistical summaries to predict the training success of users. Experiments in an industrial training task achieved a 10-20% improvement in predicting users’ training success compared to a baseline model. In addition, [273] examined skill training for searching and tracking in virtual police rooms. The authors designed virtual experiments in which officers were tasked with searching and gathering evidence for an investigation, while also considering various criminal activities. Analyses of perceptual-cognitive skills in multiple training scenarios, taking saccadic eye movements into account, indicated that training users for visual search expertise in VR is possible.
@@
Previous research has also utilized eye tracking and VR to study vision impairment and ocular examination, as vision problems lead to a perception of the world differently. Like other subfields discussed, in VR, it is possible to conduct highly controlled, programmed experiments compared to in-the-wild settings, allowing researchers to study such phenomena, which is why researchers have attempted to design ecologically valid simulations to study vision impairment issues in VR. For example, [274] simulated cataract patients’ vision in a virtual environment to better understand the perception of people with cataracts. The authors used eye tracking for these gaze-dependent symptoms because cataracts lead to special visual effects at the lens center and periphery.
@@
More related to ocular examination using eye tracking in VR, [275] transplanted the Developmental Eye Movement test to VR HMDs, which is a clinical eye test widely used to determine abnormalities in visual function and to assess ocular motor skills. Subsequently, [276] designed the King-Devick test, a standard measurement for assessing saccadic eye movement and dynamic visual acuity, using HMDs in both VR and AR. Furthermore, [277] designed a gaze-based ocular examination for visual field defects in a virtual environment using fixation and saccadic features. Compared to conventional tests that take over 30 minutes, the proposed method can be conducted in five minutes while retaining sufficient accuracy and improving reliability. Additionally, medical VR applications can extend beyond setups that require HMDs. For instance, [278] analyzed patients’ eye and head movements with homonymous visual field defects (HVFD) in a virtual driving simulation. The virtual environment featured a VR cabin, and a head-mounted eye tracker tracked participants’ eye movements. The authors’ findings confirmed their hypothesis that a particular subset of HVFD patients could enhance their viewing behaviors by increasing visual scanning, thereby improving their driving skills. Similar to earlier sections, Table VIII presents an overview of the works on human perception.
@@
TABLE VIII: Overview of the papers related to human perception. “#P” corresponds to the number of participants.
Characteristics #P Features
[239] Depth perception analysis in 3D painting 18 gaze
[240] Depth perception analysis 24 ipd, eva
[241] Continuity analysis in VR movies 49 gaze
[242] Perceptual limit analysis 37 gaze
[243] Perceptual limit analysis 27 blnk
[244] Perceptual limit analysis for redirected walking 32 blnk
[245] Perceptual limit analysis 36 sac
[246] Perceptual limit analysis 10 sac
[247] Perceptual limit analysis 13 sac
[248] Perceptual realism analysis in 3D scenes 5 gaze, pup
[249] Perception analysis in crowd walking 21 gaze, fix
[250] Perception analysis for collision avoidance 17 gaze, fix
[251] Perception analysis in psychology 30 fix
[252] Perception analysis 20 gaze, pup, blink
[253] Perceived realism analysis for virtual avatars 12 gaze
[254] Perceived realism analysis 20 gaze
[255] Perceived realism analysis for virtual avatars 18 gaze
[256] Perceived realism analysis for virtual avatars 20 gaze
[257] Face & eye synthesis for virtual avatars 18 N/A
[258] Face & eye synthesis for virtual avatars 5 N/A
[259] Face & eye synthesis for virtual avatars 7 N/A
[260] Face & eye synthesis for virtual avatars 310 N/A
[261] Face & eye synthesis for virtual avatars 3 N/A
[262] VR sickness prediction 30 gaze, pup, cvgn
[263] VR sickness prediction 30 gaze, pup, cvgn, blnk
[264] VR sickness prediction 96 saliency
[267] VR sickness mitigation 22 gaze
[269] Skill training for driving 50 gaze
[270] Skill training in education 16 pup, sac, fix
[271] Skill training for maintenance 12 fix, scanpath, saliency
[272] Skill training for sanitary apprentice 48 gaze
[273] Skill training for police room search task 54 gaze, sac, fix, entropy
[274] Vision simulation for patients with cataract 21 gaze, pup
[275] Ocular visual acuity assessment 39 eye dominance, ipd, cvgn, phr
[276] Ocular visual acuity assessment 30 sac
[277] Ocular visual field analysis 2 sac, fix
[278] Visual field analysis in driving 14 sac, fix, gaze